2x + 2y + z= 1
-x + y +2z = 3
x +2y +4z =0
Help! Solve 3 step linear equation

Answer by TimothyLamb(4379)

---
2x + 2y + z= 1
-x + y +2z = 3
x +2y +4z =0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -2
y = 3
z = -1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php